OpenStudy (anonymous):

Jamie bought 15 apples. He ate 3 apples. What percentage of the apples did Jamie eat?

OpenStudy (godlovesme):

3/15 = x/ 100 cross multiply and solve for x make sense? @smartkid2000

OpenStudy (anonymous):

?

OpenStudy (anonymous):

I WILL MEDAL IF YOU GIVE ME ANSWER PLS

OpenStudy (godlovesme):

Sorry I can't give direct answer 3/15= x/ 100 15x = 300 X=20% there u go hope that helped C: @smartkid2000